OpenStudy (anonymous):

PLEASE HELP THANK YOU. How to solve for m 3^15+3^15+3^15=3^m

OpenStudy (j-cross18):

Are you in Geometry or algebra 2??

OpenStudy (anonymous):

algebra :)

OpenStudy (j-cross18):

one or two because i might be able to solve this...

OpenStudy (anonymous):

I'm in Algebra Honors, 8th Grade.

OpenStudy (j-cross18):

OK i got 43,046,721=3^m Now to simply you do what??? Do you know???

OpenStudy (j-cross18):

@megansri29

OpenStudy (anonymous):

notice you can factor out a \[3^{15}\] and now you have \[3^{15}(1+1+1) = 3^m\] \[3^{15}(3) = 3^m\] \[3^{16} = 3^m\] solve for my from here

OpenStudy (anonymous):

m*

OpenStudy (anonymous):

oh alright! thank you so much!

OpenStudy (anonymous):

your welcome glad to help

OpenStudy (anonymous):

thanks again :D

OpenStudy (j-cross18):

you are welcome so if you want you can give me a medal for the effort plz thanks ^.^

OpenStudy (j-cross18):

Are you @megansri29 ?????

OpenStudy (j-cross18):

i think you gave the medal to the wrong person...